Please I need help!!
5-(9-6x)=2x-2
7-(5-8x)=4x+3
if 7x+3=5x-7,evaluate 3x-2

Answer by stanbon(75887)

5-(9-6x)=2x-2
5-9+6x = 2x-2
6x-4 = 2x-2
4x = 6
x = 3/2
-----------------
7-(5-8x)=4x+3
7-5+8x = 4x+3
2+8x = 4x+3
4x = 1
x = 1/4
-----------------
if 7x+3=5x-7,evaluate 3x-2
2x = -10
x = -5
-----
3x = -15
3x-2 = -17
=====================
Cheers,
Stan H.
=====================